OpenStudy (anonymous):

I do not understand someone please explain! Let n be the middle number of three consecutive even integers. Write an expression for the sum of these integers.

OpenStudy (unklerhaukus):

consecutive numbers are numbers that are all in a row; if the first of four consecutive numbers is x the numbers will be x, x+1, x+2, x+3

OpenStudy (unklerhaukus):

In your case, you only need three consecutive numbers, and the middle one is n

OpenStudy (anonymous):

I think I get it now. Would it be sum=3n?

OpenStudy (solomonzelman):

if the sum of these 3 consec. integers is equal to certain number C and you want to find the numbers, I would write \(\large\color{black}{ \displaystyle (x-1)+(x)+(x+1) =C }\)

OpenStudy (solomonzelman):

making x the middle nuber

OpenStudy (unklerhaukus):

yes!, the numbers are n-1, n, n+1 so their sum is (n-1) + n + (n+1) = 3n

OpenStudy (anonymous):

But it 3 consecutive EVEN integers so wouldn't 2 be added each time?

OpenStudy (solomonzelman):

yes, the 1's cancel out

OpenStudy (solomonzelman):

and so it would be with n \(\large\color{black}{ \displaystyle (n-2)+(n)+(n+2) =3n }\)

OpenStudy (solomonzelman):

still 3n

OpenStudy (anonymous):

Thank you!

OpenStudy (unklerhaukus):

Dammit, i didn't read the question correctly, consecutive even numbers are x, x+2, x+4, ... so (n-2) + n + (n+2) = same result

OpenStudy (solomonzelman):

I miss-read it at first too

OpenStudy (solomonzelman):

but, for me it is typical

OpenStudy (unklerhaukus):

(me too)